OpenStudy (anonymous):

Deriving Physics Equations? My textbook is not clear and I need help with how they went from step 3 to 4.

OpenStudy (anonymous):

Step 1: \[X = v _{i}t\] Step 2: \[Y = \frac{ 1 }{ 2 }t(v _{f}-v _{i})\] Step 3: \[X + Y = v _{i}t+(\frac{ v _{f} t }{ 2 }-\frac{ v _{i} t }{ 2 })\] Step 4: \[d = (\frac{ v _{f}+v _{i} }{ 2 })t\]

jimthompson5910 (jim_thompson5910):

I'm starting with step 3 \[X + Y = v _{i}t+(\frac{ v _{f} t }{ 2 }-\frac{ v _{i} t }{ 2 })\] \[X + Y = \frac{2v _{i}t}{2}+(\frac{ v _{f} t }{ 2 }-\frac{ v _{i} t }{ 2 })\] \[X + Y = \frac{2v _{i}t}{2}+\frac{ v _{f} t }{ 2 }-\frac{ v _{i} t }{ 2 }\] \[X + Y = \frac{2v _{i}t}{2}-\frac{ v _{i} t }{ 2 }+\frac{ v _{f} t }{ 2 }\] \[X + Y = \frac{2v _{i}t-v _{i} t}{2}+\frac{ v _{f} t }{ 2 }\] \[X + Y = \frac{v _{i} t}{2}+\frac{ v _{f} t }{ 2 }\] \[X + Y = \frac{v _{i} t+ v _{f} t }{ 2 }\] \[X + Y = \frac{(v _{i} + v _{f}) t }{ 2 }\] \[X + Y = \left(\frac{v _{i} + v _{f} }{ 2 }\right)t\] \[d = \left(\frac{v _{i} + v _{f} }{ 2 }\right)t\]

OpenStudy (anonymous):

thanks @jim_thompson5910